If the Moon were twice as massive as it is now, and it has stayed in the same orbital radius about the
Earth as it has now, its new orbital period (in terms of its current orbital period T ) would be,
A) T (B) 2T
(C) T/2
(D) T/4

Grade:10

1 Answers

vikas askiitian expert
13 years ago

orbital velocity = V =(GM/R)1/2

 finally mass is doubled so  orbital velocity becomes V1 = (2GM/R)1/2

T = 2piR/V              ( V is orbital velocity & R is orbital radius , T is time period)

T1/T = V/V1

 T1 = T/sqrt2